Question 1 Bridgeport provides environmentally friendly lawn services for homeowners. Its operating costs are as follows. Depreciation $1,900 per month Advertising $300 per month Insurance $3,080 per month Weed and feed materials $13 per lawn Direct labor $16 per lawn Fuel $3 per lawn Bridgeport charges $80 per treatment for the average single-family lawn. Collapse question part (a) This grade has been saved to Gradebook. Your answer has been saved and sent for grading. See Gradebook for score details. Determine the company’s break-even point in number of lawns serviced per month. Break-even point lawns LINK TO TEXT LINK TO TEXT Attempts: 1 of 1 used Collapse question part (b) This grade has been saved to Gradebook. Your answer has been saved and sent for grading. See Gradebook for score details. Determine the company’s break-even point in dollars. (Round answer to 0 decimal places, e.g. 1,225.) Break-even point $
Answer #1
Total fixed cost 5280 =1900+300+3080
Variable cost per lawn 32 =13+16+3
Contribution margin per lawn 48 =80-32
Contribution margin ratio 60% =48/80
a
Fixed costs 5280
Divide by Contribution margin per lawn 48
Break-even point lawns 110
b
Fixed costs 5280
Divide by Contribution margin ratio 60%
Break-even point $ 8800